# RASPNet: A Benchmark Dataset for Radar Adaptive Signal Processing Applications

### v2

@@ G.4 Challenging Scenarios
[Uncaptioned image] Platform Aim Location (lat, lon): $(45.5953^{\circ},-84.9085^{\circ})$ Platform Position (lat, lon): $(45.7183^{\circ},-87.7585^{\circ})$ Folder: /CLUTTER/num74_lat45.7183_lon-87.7585/ Description: Scenario $q_{i}=51\ (i=74)$ in RASPNet is near the Straits of Mackinac in Michigan. The radar platform is positioned at latitude $45.7183^{\circ}$ N and longitude $-87.7585^{\circ}$ W. It aims towards a location at latitude $45.5953^{\circ}$ N and longitude $-84.9085^{\circ}$ W. The radar is positioned near the coastline, the aim direction is inland, and the terrain has moderate variability.

### v3

@@ G.4 Challenging Scenarios
[Uncaptioned image] Platform Aim Location (lat, lon): $(45.5953^{\circ},-84.9085^{\circ})$ Platform Position (lat, lon): $(45.7183^{\circ},-87.7585^{\circ})$ Folder: /CLUTTER/num74_lat45.7183_lon-87.7585/ Description: Scenario $q_{i}=51\ (i=74)$ in RASPNet is near the Straits of Mackinac in Michigan. The radar platform is positioned at latitude $45.7183^{\circ}$ N and longitude $87.7585^{\circ}$ W. It aims towards a location at latitude $45.5953^{\circ}$ N and longitude $84.9085^{\circ}$ W. The radar is positioned near the coastline, the aim direction is inland, and the terrain has moderate variability.